OpenStudy (anonymous):

4xy^0/z^-3 A. z^-3 B. 1/z^-3 C. 4z^3 D. 4xz^3 please help

jimthompson5910 (jim_thompson5910):

Hints: y^0 = 1 this is true for any variable really and x^(-k) = 1/(x^k) 1/(x^(-k)) = x^k